The sum of the digits of a two-digit number is 9. The value of the number is 12 times the tens digit. Find the number
Answer by ankor@dixie-net.com(12704)

let x = 1st digit (tens), y = 2nd digit (ones)
:
The sum of the digits of a two-digit number is 9.
x + y = 9
or
x = 9-y, will us this for substitution

The value of the number is 12 times the tens digit. Find the number
10x + y = 12(x)
y = 12x - 10x
y = 2x
:
Substitute (9-y) for x
y = 2(9-y)
y = 18 - 2y
y + 2y = 18
3y = 18
y = 6
:
x = 3,
:
The number is 36
:
Check: 36 is 12 times the 10's digit which is 3